OpenStudy (chycora):

1. When Elyse moved last weekend, the circular glass top for an end table broke. To replace it, she bought a circular glass top with a diameter of 24 inches. What is the approximate area of the circular glass top Elyse bought? (Use 3.14 for .) 452.16 in2 150.72 in2 75.36 in2 1,808.64 in2

OpenStudy (asad786):

i had this on study island

OpenStudy (chycora):

kool u go 2 k12

OpenStudy (asad786):

yes

OpenStudy (asad786):

im not sure i forgot this question

OpenStudy (asad786):

but 3.14 times 24 try that

OpenStudy (asad786):

wait don't answer yet

OpenStudy (asad786):

@StudyGurl14

OpenStudy (asad786):

@happydeedee

OpenStudy (asad786):

@iGreen

OpenStudy (studygurl14):

Formula for area of circle: \(\large r^2\pi\) r = radius Remember that diameter = 2(radius_

OpenStudy (asad786):

so its ridius times 2 right

OpenStudy (asad786):

sorry couldn't spell that right

OpenStudy (asad786):

ok so 24 times 3.14 should be the answer

OpenStudy (chycora):

k so 75.36

OpenStudy (asad786):

tag studygurl14

OpenStudy (asad786):

yes

OpenStudy (studygurl14):

@Asad786 You're wrong. It's not 24 times 3.14

OpenStudy (studygurl14):

24 is the diameter, not the radius. You want the radius.

OpenStudy (asad786):

oh so 12 times 3.14

OpenStudy (studygurl14):

No. Remember the formula is \(\large r^2\pi\)

OpenStudy (asad786):

r times 2 right

OpenStudy (chycora):

it is A

OpenStudy (asad786):

no i remember

OpenStudy (studygurl14):

No, r squared.

OpenStudy (studygurl14):

Yes @chycora Nice job. A is correct

OpenStudy (asad786):

tho for future reference can someone teach me

OpenStudy (studygurl14):

Yeah @Asad786 . Open a new thread with it and tag me

OpenStudy (asad786):

k

OpenStudy (chycora):

yea thanx guys lol

OpenStudy (studygurl14):

you're welocme @chycora